p(2ap,ap^2) is any point on the parabolax^2 = 4ay. The normal at p meets the y-axis at g. N is a point such that p is the midpoint of GN. Show that the equation of the locus of N is a parabola and find its vertex and focal length.

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
slope of tangent to given parabola= at g is 
m1=(dy/dx) at g= 2/8= ¼
also slope at N, m2= (dy/dx) at N= 1
as we found that y^2=4x
here a=1, therefore it is a parabola. [proved]
now,vertex=(0,0)
and focal length=a=1 units [ans]
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)